OpenStudy (anonymous):

write the expression (-216)^4/3 in radical form.

OpenStudy (anonymous):

the 4 is a power and the 3 is an index\[\sqrt[3]{-216^{4}}\]

OpenStudy (anonymous):

do you know how to evaluate the radical form?

OpenStudy (anonymous):

\[\sqrt[3]{(-216)^4}=\sqrt[3]{2176782336}=1296 \]

OpenStudy (anonymous):

\[\sqrt[3]{((-2)^{3}(3)^{3})^{4}}\] \[so you are lefts with (-2 \times 3) raised \to the 4th power\]

OpenStudy (anonymous):

the threes under the radical and the 3 index cancel. sorry about earlier post. having trouble with the equation button

OpenStudy (anonymous):

its fine.